# On the Mode-Seeking Properties of Langevin Dynamics

### v2

@@ Appendix C Convergence Analysis of Chained Langevin Dynamics
$\text{TV}\left(\hat{P}\left({\mathbf{x}}^{[q]}\right),P\left({\mathbf{x}}^{[q]}\right)\right)$
$=\frac{1}{2}\int\left\lvert\hat{P}\left({\mathbf{x}}^{[q]}\right)-P\left({\mathbf{x}}^{[q]}\right)\right\rvert{\,\mathrm{d}}{\mathbf{x}}^{[q]}$
$=\frac{1}{2}\int\left\lvert\hat{P}\left({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{[q-1]}\right)\hat{P}\left({\mathbf{x}}^{[q-1]}\right)-P\left({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{[q-1]}\right)P\left({\mathbf{x}}^{[q-1]}\right)\right\rvert{\,\mathrm{d}}{\mathbf{x}}^{[q]}$
$\leq\frac{1}{2}\int\left\lvert\hat{P}\left({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{[q-1]}\right)\hat{P}\left({\mathbf{x}}^{[q-1]}\right)-\hat{P}\left({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{[q-1]}\right)P\left({\mathbf{x}}^{[q-1]}\right)\right\rvert{\,\mathrm{d}}{\mathbf{x}}^{[q]}$
$\quad\quad+\frac{1}{2}\int\left\lvert\hat{P}\left({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{[q-1]}\right)P\left({\mathbf{x}}^{[q-1]}\right)-P\left({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{[q-1]}\right)P\left({\mathbf{x}}^{[q-1]}\right)\right\rvert{\,\mathrm{d}}{\mathbf{x}}^{[q]}$
$=\frac{1}{2}\int\hat{P}\left({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{[q-1]}\right){\,\mathrm{d}}{\mathbf{x}}^{(q)}\int\left\lvert\hat{P}\left({\mathbf{x}}^{[q-1]}\right)-P\left({\mathbf{x}}^{[q-1]}\right)\right\rvert{\,\mathrm{d}}{\mathbf{x}}^{[q-1]}$
$\quad\quad+\frac{1}{2}\int\left\lvert\hat{P}\left({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{[q-1]}\right)-P\left({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{[q-1]}\right)\right\rvert{\,\mathrm{d}}{\mathbf{x}}^{(q)}\int P\left({\mathbf{x}}^{[q-1]}\right){\,\mathrm{d}}{\mathbf{x}}^{[q-1]}$
$=\text{TV}\left(\hat{P}\left({\mathbf{x}}^{[q-1]}\right),P\left({\mathbf{x}}^{[q-1]}\right)\right)+\text{TV}\left(\hat{P}\left({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{[q-1]}\right),P\left({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{[q-1]}\right)\right)$
$\leq\text{TV}\left(\hat{P}\left({\mathbf{x}}^{[q-1]}\right),P\left({\mathbf{x}}^{[q-1]}\right)\right)+\varepsilon\cdot\frac{Q}{d}.$
@@
$\text{TV}\left(\hat{P}({\mathbf{x}}),P({\mathbf{x}})\right)$ $=\sum_{q=1}^{d/Q}\left(\text{TV}\left(\hat{P}\left({\mathbf{x}}^{[q]}\right),P\left({\mathbf{x}}^{[q]}\right)\right)-\text{TV}\left(\hat{P}\left({\mathbf{x}}^{[q-1]}\right),P\left({\mathbf{x}}^{[q-1]}\right)\right)\right)$
@@ D.1 Synthetic Gaussian Mixture Model
Since the distribution is given, we assume that the sampling algorithms have access to the ground-truth score function. We set the batch size as 1000 and patch size $Q=10$ for chained Langevin dynamics. We use $T\in\left\{{10^{3},10^{4},10^{5},10^{6}}\right\}$ iterations for vanilla and chained Langevin dynamics. A sample ${\mathbf{x}}$ is clustered in mode 1 if it satisfies $\left\|{{\mathbf{x}}-{\bm{\mu}}_{1}}\right\|^{2}\leq 5d$ and $\left\|{{\mathbf{x}}-{\bm{\mu}}_{1}}\right\|^{2}\leq\left\|{{\mathbf{x}}-{\bm{\mu}}_{2}}\right\|^{2}$; in mode 2 if $\left\|{{\mathbf{x}}-{\bm{\mu}}_{2}}\right\|^{2}\leq 5d$ and $\left\|{{\mathbf{x}}-{\bm{\mu}}_{1}}\right\|^{2}>\left\|{{\mathbf{x}}-{\bm{\mu}}_{2}}\right\|^{2}$; and in mode 0 otherwise. The initial samples are i.i.d. chosen from $P^{(0)}$, $P^{(1)}$, or $P^{(2)}$, and the results are presented in Figures 2, 5, and 6 respectively. The two subfigures above the dashed line illustrate the samples from the initial distribution and target distribution, and the subfigures below the dashed line are the samples generated by different algorithms. Furthermore, in Figures 7, 8 and 9 we demonstrate the effect of different values of $Q\in\{1,4,10,20,50\}$ on the mode-seeking tendencies of Chained-LD. We can observe that for dimension $d=100$, a moderate patch size $Q\in\{1,4,10\}$ has similar performance, a large patch size $Q=20$ needs more steps to find the other two modes, while an overly-large patch size $Q=50$ almost cannot find other modes.
@@ D.3 Image Datasets
Our implementation and hyperparameter selection are based on [4] and [26]. During training, we i.i.d. randomly flip an image with probability 0.5 to construct the two modes (i.e., original and flipped images). All models are optimized by Adam with learning rate 0.001 and batch size 128 for a total of 200000 training steps, and we use the model at the last iteration to generate the samples. We perform experiments on MNIST [41] (CC BY-SA 3.0 License) and Fashion-MNIST [42] (MIT License) datasets and we set the patch size as $Q=14$.
@@
To generate samples, we use $T\in\left\{{10000,30000,100000}\right\}$ iterations for annealed Langevin dynamics (ALD) and Chained-ALD. The initial samples are chosen as either original or flipped images from the dataset, and the results for MNIST and Fashion-MNIST datasets are presented in Figures 11, 3, 4, and 12 respectively. The two subfigures above the dashed line illustrate the samples from the initial distribution and target distribution, and the subfigures below the dashed line are the samples generated by different algorithms.
All experiments were run with one RTX3090 GPU. It is worth noting that the training and inference time of chained Langevin dynamics using LSTM is considerably faster than vanilla/annealed Langevin dynamics using RefineNet. For a course of 200000 training steps on MNIST/Fashion-MNIST, due to the different network architectures, LSTM takes around 2.3 hours while RefineNet takes around 9.2 hours. Concerning image generation, chained Langevin dynamics is significantly faster than vanilla/annealed Langevin dynamics since every iteration of chained Langevin dynamics only updates a patch of constant size, while every iteration of vanilla/annealed Langevin dynamics requires computing all coordinates of the sample. One iteration of chained Langevin dynamics using LSTM takes around 1.97 ms, while one iteration of vanilla/annealed Langevin dynamics using RefineNet takes around 43.7 ms.

### v3

@@ Appendix C Proof of Theorem 4
$\text{TV}\left(\widehat{P}\left({\mathbf{x}}^{[q]}\right),P\left({\mathbf{x}}^{[q]}\right)\right)$
$=\frac{1}{2}\int\left\lvert\widehat{P}\left({\mathbf{x}}^{[q]}\right)-P\left({\mathbf{x}}^{[q]}\right)\right\rvert{\,\mathrm{d}}{\mathbf{x}}^{[q]}$
$=\frac{1}{2}\int\left\lvert\widehat{P}\left({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{[q-1]}\right)\widehat{P}\left({\mathbf{x}}^{[q-1]}\right)-P\left({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{[q-1]}\right)P\left({\mathbf{x}}^{[q-1]}\right)\right\rvert{\,\mathrm{d}}{\mathbf{x}}^{[q]}$
$\leq\frac{1}{2}\int\left\lvert\widehat{P}\left({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{[q-1]}\right)\widehat{P}\left({\mathbf{x}}^{[q-1]}\right)-\widehat{P}\left({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{[q-1]}\right)P\left({\mathbf{x}}^{[q-1]}\right)\right\rvert{\,\mathrm{d}}{\mathbf{x}}^{[q]}$
$\quad\quad+\frac{1}{2}\int\left\lvert\widehat{P}\left({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{[q-1]}\right)P\left({\mathbf{x}}^{[q-1]}\right)-P\left({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{[q-1]}\right)P\left({\mathbf{x}}^{[q-1]}\right)\right\rvert{\,\mathrm{d}}{\mathbf{x}}^{[q]}$
$=\frac{1}{2}\int\widehat{P}\left({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{[q-1]}\right){\,\mathrm{d}}{\mathbf{x}}^{(q)}\int\left\lvert\widehat{P}\left({\mathbf{x}}^{[q-1]}\right)-P\left({\mathbf{x}}^{[q-1]}\right)\right\rvert{\,\mathrm{d}}{\mathbf{x}}^{[q-1]}$
$\quad\quad+\frac{1}{2}\int\left\lvert\widehat{P}\left({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{[q-1]}\right)-P\left({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{[q-1]}\right)\right\rvert{\,\mathrm{d}}{\mathbf{x}}^{(q)}\int P\left({\mathbf{x}}^{[q-1]}\right){\,\mathrm{d}}{\mathbf{x}}^{[q-1]}$
$=\text{TV}\left(\widehat{P}\left({\mathbf{x}}^{[q-1]}\right),P\left({\mathbf{x}}^{[q-1]}\right)\right)+\text{TV}\left(\widehat{P}\left({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{[q-1]}\right),P\left({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{[q-1]}\right)\right)$
$\leq\text{TV}\left(\widehat{P}\left({\mathbf{x}}^{[q-1]}\right),P\left({\mathbf{x}}^{[q-1]}\right)\right)+\varepsilon\cdot\frac{Q}{d}.$
@@
$\text{TV}\left(\widehat{P}({\mathbf{x}}),P({\mathbf{x}})\right)$ $=\sum_{q=1}^{d/Q}\left(\text{TV}\left(\widehat{P}\left({\mathbf{x}}^{[q]}\right),P\left({\mathbf{x}}^{[q]}\right)\right)-\text{TV}\left(\widehat{P}\left({\mathbf{x}}^{[q-1]}\right),P\left({\mathbf{x}}^{[q-1]}\right)\right)\right)$
@@
Finally, upon combining Propositions 2 and 3, we finish the proof of Theorem 4.
@@ D.1 Synthetic Gaussian Mixture Model
Since the distribution is given, we assume that the sampling algorithms have access to the ground-truth score function. We set the batch size as 1000 and patch size $Q=10$ for chained Langevin dynamics. We use $T\in\left\{{10^{3},10^{4},10^{5},10^{6}}\right\}$ iterations for vanilla and chained Langevin dynamics. A sample ${\mathbf{x}}$ is clustered in mode 1 if it satisfies $\left\|{{\mathbf{x}}-{\bm{\mu}}_{1}}\right\|^{2}\leq 5d$ and $\left\|{{\mathbf{x}}-{\bm{\mu}}_{1}}\right\|^{2}\leq\left\|{{\mathbf{x}}-{\bm{\mu}}_{2}}\right\|^{2}$; in mode 2 if $\left\|{{\mathbf{x}}-{\bm{\mu}}_{2}}\right\|^{2}\leq 5d$ and $\left\|{{\mathbf{x}}-{\bm{\mu}}_{1}}\right\|^{2}>\left\|{{\mathbf{x}}-{\bm{\mu}}_{2}}\right\|^{2}$; and in mode 0 otherwise. The initial samples are i.i.d. chosen from $P^{(0)}$, $P^{(1)}$, or $P^{(2)}$, and the results are presented in Figures 3, 5, and 6 respectively. The two subfigures above the dashed line illustrate the samples from the initial distribution and target distribution, and the subfigures below the dashed line are the samples generated by different algorithms. Furthermore, in Figures 7, 8 and 9 we demonstrate the effect of different values of $Q\in\{1,4,10,20,50\}$ on the convergence of Chained-LD. We can observe that for dimension $d=100$, a moderate patch size $Q\in\{1,4,10\}$ has similar performance, a large patch size $Q=20$ needs more steps to find the other two modes, while an overly-large patch size $Q=50$ almost cannot find other modes.
We further numerically evaluate the performance of LD and Chained-LD in other Gaussian mixture models. We consider an in-between mode $P^{(0)}={\mathcal{N}}(\bm{0}_{d},10{\bm{I}}_{d})$ in dimension $d=100$ with weight $w_{0}=0.01$, and the other modes have the same weight, the same covariance matrix but different mean, i.e., $P^{(i)}={\mathcal{N}}({\bm{\mu}}_{i},0.1{\bm{I}}_{d})$ and $w_{i}=0.99/k$. The first two coordinates of ${\bm{\mu}}_{i}$ are chosen as shown in Figures 11 and 12, and the other coordinates of ${\bm{\mu}}_{i}$ are set to be 0. The numerical results in Figures 11 and 12 are consistent with our previous analysis.
@@ D.3 Image Datasets
Our implementation and hyperparameter selection are based on [6] and [26]. During training, we i.i.d. randomly flip an image with probability 0.5 to construct the two modes (i.e., original and flipped images). All models are optimized by Adam with learning rate 0.001 and batch size 128 for a total of 200000 training steps, and we use the model at the last iteration to generate the samples. We perform experiments on MNIST [28] (CC BY-SA 3.0 License) and Fashion-MNIST [29] (MIT License) datasets and we set the patch size as $Q=14$. All experiments were run with one RTX3090 GPU.
@@
To generate samples, we use $T\in\left\{{10000,30000,100000}\right\}$ iterations for annealed Langevin dynamics (ALD) and Chained-ALD. The initial samples are chosen as either original or flipped images from the dataset, and the results for MNIST and Fashion-MNIST datasets are presented in Figures 13, 4, 14, and 15 respectively. The two subfigures above the dashed line illustrate the samples from the initial distribution and target distribution, and the subfigures below the dashed line are the samples generated by different algorithms.
@@
Figure 11: Samples from a mixture of 9 Gaussian modes (including an in-between mode $P^{(0)}={\mathcal{N}}(\bm{0}_{d},10{\bm{I}}_{d})$) generated by vanilla Langevin dynamics (VLD) and chained vanilla Langevin dynamics (Chained-VLD) with patch size $Q=1$. Two axes are the first 2 coordinates of the samples. The samples are initialized in ${\mathcal{N}}(\bm{0}_{100},{\bm{I}}_{100})$.
Figure 12: Samples from a mixture of 10 Gaussian modes (including an in-between mode $P^{(0)}={\mathcal{N}}(\bm{0}_{d},10{\bm{I}}_{d})$) generated by vanilla Langevin dynamics (VLD) and chained vanilla Langevin dynamics (Chained-VLD) with patch size $Q=1$. Two axes are the first 2 coordinates of the samples. The samples are initialized in ${\mathcal{N}}(\bm{0}_{100},{\bm{I}}_{100})$.